OpenStudy (anonymous):

Help...Find the point slope form of an equation of the line through the two ordered pairs (7,20) (16,72)

jigglypuff314 (jigglypuff314):

\[slope = m = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] when given the points (x1, y1) and (x2, y2) once you found the slope, plug in what you know into the point slope form y - y1 = m(x - x1)

OpenStudy (anonymous):

Thanks